# SoK: A Literature and Engineering Review of Regular Expression Denial of Service (ReDoS)

### v3

@@ 3.1.1. Detection
Dynamic Analysis (4 studies) Dynamic approaches (Shen et al., 2018; McLaughlin et al., 2022; Barlas et al., 2022; Petsios et al., 2017) to detecting ReDoS vulnerabilities involve executing regular expressions with various inputs and analyzing their runtime behavior to identify potential issues. For instance, Shen et al. (Shen et al., 2018) developed ReScue, a tool that uses an evolutionary genetic algorithm to generate time-consuming strings capable of triggering ReDoS vulnerabilities. Similarly, Barlas et al. (Barlas et al., 2022) and McLaughlin et al. (McLaughlin et al., 2022) utilized dynamic analysis with fuzzing techniques, while Petsios et al. (Petsios et al., 2017) introduced SlowFuzz, which uses automated fuzzing to find inputs causing worst-case algorithmic behavior.
Hybrid Approaches (6 studies) Existing static analysis methods for detecting ReDoS vulnerabilities often face a trade-off between precision and recall. For example, Davis et al. (Davis et al., 2018a) and Shen et al. (Shen et al., 2018) reported low F-1 scores of 44.94% and 3.57%, respectively (Li et al., 2021). To address these limitations, dynamic validation is frequently employed alongside static analysis, resulting in hybrid approaches. Tools such as NFAA (Wüstholz et al., 2017), Revealer (Liu et al., 2021), ReDoSHunter (Li et al., 2021), Badger (Noller et al., 2018), and RENGAR (Wang et al., 2023) combine static and dynamic methods to mitigate false positives. For instance, hybrid approaches like ReDoSHunter integrate static and dynamic techniques to improve detection accuracy, while tools such as NFAA refine dynamic testing with insights from static analysis (Wüstholz et al., 2017). These methods effectively balance the strengths of both static and dynamic detection, reducing the likelihood of inaccurate results.
Machine Learning Approaches (1 studies) Machine learning techniques are also gaining traction in ReDoS detection. Unlike hybrid methods, which explicitly integrate static and dynamic analysis, machine learning approaches aim to predict vulnerabilities and detect anomalies based on patterns in data. For example, Demoulin et al. (Demoulin et al., 2019) introduced a semi-supervised learning model to analyze resource utilization and identify potential vulnerabilities. This method triggers immediate alerts and provides mitigation strategies by detecting anomalies in resource consumption. While machine learning approaches offer promise, they face challenges such as reliance on labeled data, vulnerability to adversarial inputs, and difficulty in addressing context-specific patterns that may not generalize well across diverse regex use cases.
While dynamic and hybrid techniques are more precise, they tend to be slow. Parolini et al. (Parolini and Miné, 2023) show that ReScue (Shen et al., 2018) and Rexploiter (Wüstholz et al., 2017) are 2055x and 57x slower than RXXR2 (Rathnayake and Thielecke, 2017). Thus, despite their strenghts, these methods exhibit slow execution time and incomplete coverage.
@@ 3.1.2. Prevention
Better Regex Matching Algorithms (3 studies) Researchers have proposed developing new regex matching engines to address ReDoS vulnerabilities at a fundamental level. Davis et al. (Davis et al., 2021; Davis, 2019) introduced memoization-based optimization to speed up regex matching by caching intermediate results, significantly reducing redundant computations. Turoňová et al. (Turoňová et al., 2020, 2022) proposed novel counting automata matching algorithms that further optimize counter management during matching to efficiently handle regexes with bounded repetitions. These new engines often employ finite-state machine-based approaches instead of backtracking, inherently avoiding the exponential time complexity that leads to ReDoS vulnerabilities. Additionally, some methods impose restrictions on regex features to ensure faster and safer processing.
Automatic Repair (4 studies) Automatic repair techniques aim to proactively address vulnerabilities in regular expressions to mitigate ReDoS attacks. Several studies have explored programmatic frameworks to detect and repair problematic regex patterns. For example, Li et al. (Li et al., 2020) developed FlashRegex, a programming-by-example (PBE) tool that repairs or synthesizes regex from provided matching examples. However, FlashRegex struggles with complex regex features, such as lookarounds and backreferences (Li et al., 2022). To address these limitations, RegexScalpel (Li et al., 2022) employs a localize-and-fix approach, which identifies fine-grained vulnerabilities and applies predefined repair strategies to improve recall and precision. In a complementary effort, Hassan et al. (Hassan et al., 2023) focus on identifying specific regex structures or “anti-patterns” that are prone to ReDoS vulnerabilities. Their method analyzes regex patterns to uncover and repair potential vulnerabilities early in the development process, avoiding the need for runtime execution. This approach complements programmatic frameworks by systematically identifying and addressing problematic constructs.
Other tools, such as Chida et al. (Chida and Terauchi, 2022), also leverage PBE algorithms but face challenges related to dependency on user-provided examples, which can result in incomplete repairs (Li et al., 2021; Wang et al., 2023). These methods, while effective for certain regex patterns, often struggle with advanced or highly customized regex constructs, underscoring the need for more comprehensive solutions capable of addressing complex vulnerabilities. Moreover, improving the generalization and robustness of automatic repair frameworks remains a key area of ongoing research.
@@ 3.1.3. Mitigation
Anomaly Detection (2 studies) Anomaly detection techniques are central to mitigation strategies, leveraging real-time monitoring to identify unusual patterns indicative of ReDoS attacks. Machine learning (ML)-based approaches have proven effective in this domain (Bai et al., 2021; Tandon et al., 2023). For example, Tandon et al. (Tandon et al., 2023) employ the Elliptic Envelope model to analyze runtime features such as request time, memory usage, and CPU cycles, enabling the system to detect deviations from normal behavior. Bai et al. (Bai et al., 2021) introduce a feedback loop mechanism that continuously adjusts defensive measures based on real-time performance metrics, providing robust protection against zero-day ReDoS attacks.
@@
Limiting Resource Consumption (1 study) To mitigate ReDoS attacks, Davis et al. (Davis et al., 2018b) proposed using first-class timeouts as a defense against Event Handler Poisoning (EHP) in Node.js. Their prototype, Node.cure, effectively limits processing time with minimal performance overhead. Similarly, runtime environments like PHP, Perl, and .NET have adopted resource-limiting mechanisms to curb backtracking and excessive processing. In PHP, the pcre.backtrack_limit and pcre.recursion_limit directives control the number of backtracks and recursion depth during regex matching (PHP Documentation Team, 2025). Perl’s PCRE2 library allows setting evaluation limits via the eval function, with a default recursion depth limit of 10M. In .NET, the RegexOptions.MatchTimeout property enables developers to specify a maximum regex processing duration. While these strategies effectively cap resources to reduce ReDoS risks, they are often too coarse-grained, making it challenging to balance performance and avoid false positives in complex regex patterns or large datasets.
@@ 3.1.4. Other Related Work
Dataset Studies Bhuiyan et al. (Bhuiyan et al., 2023) introduced SecBenchJS, a benchmark dataset designed to evaluate the security of JavaScript applications, including regex vulnerabilities. Such datasets enable standardized testing and improve ReDoS detection tools.
Empirical Studies Staicu et al. (Staicu and Pradel, 2018) analyzed 2,846 popular websites, finding 339 vulnerable to ReDoS attacks, highlighting the risks of server-side JavaScript. Davis et al. (Davis et al., 2019) studied regex portability across eight programming languages, uncovering semantic and performance differences in 25% of regexes. These studies underscore the need for safer regex practices and consistent regex engine implementations.
LLM-Based Studies Siddiq et al. (Siddiq et al., 2024b) proposed Re(gEx—DoS)Eval, a framework for evaluating LLM-generated regexes using metrics like pass@k and vulnerable@k, demonstrating its use on T5, Phi, and GPT-3. In another study, Siddiq et al. (Siddiq et al., 2024a) categorized vulnerable regex patterns and analyzed developer discussions, finding that GPT-3.5 often generates regexes with polynomial vulnerabilities.
@@ 3.2. Threat Models
Table 2. Overview of threat models considered in the reviewed studies. Most works consider regexes in isolation or only within a library; only 7/35 of these works consider reachability at the application level. Beyond this, ReDoS research makes strong assumptions about the attacker’s capabilities (or assumes a poorly defended system).
Paper Code Fragment Library Application Arbitrary Number of Requests Su et al. (Su et al., 2024) 100kB Davis et al. (Davis et al., 2017) Turoňová et al. (Turoňová et al., 2022) 1.5K-9K Liu et al. (Liu et al., 2021) 128 Staicu et al. (Staicu and Pradel, 2018) 82K Davis et al. (Davis et al., 2018a) 100K-1M Siddiq et al. (Siddiq et al., 2024a) Shen et al. (Shen et al., 2018) 128 Barlas et al. (Barlas et al., 2022) Siddiq et al. (Siddiq et al., 2024b) Davis et al. (Davis et al., 2019) Kluban et al. (Kluban et al., 2022) Bhuiyan et al. (Bhuiyan et al., 2023) Noller et al. (Noller et al., 2018) Wang et al. (Wang et al., 2023) Rathnayake et al. (Rathnayake and Thielecke, 2017) McLaughlin et al. (McLaughlin et al., 2022) 1M Li et al. (Li et al., 2021) Parolini et al. (Parolini and Miné, 2023) 128 Kirrage et al. (Kirrage et al., 2013) Wüstholz et al. (Wüstholz et al., 2017) 140K Petsios et al. (Petsios et al., 2017) Kluban et al. (Kluban et al., 2024) Demoulin et al. (Demoulin et al., 2019) Chida et al. (Chida and Terauchi, 2022) Bai et al. (Bai et al., 2021) 50K Li et al. (Li et al., 2022) 1M Davis et al. (Davis, 2019) Li et al. (Li et al., 2020) Hassan et al. (Hassan et al., 2023) Davis et al. (Davis et al., 2018b) Davis et al. (Davis et al., 2021) 10K-100K Tandon et al. (Tandon et al., 2023) Weideman et al. (Weideman et al., 2016) Turoňová et al. (Turoňová et al., 2020) 500K-50M
A realistic threat model is essential in any security study, as it provides a framework for assessing the potential findings. That is, it enables the identification of system’s weaknesses, i.e., vulnerabilities, that adversaries could exploit in real-world environments. A strong threat model might lead to unrealistic findings, assuming too powerful attackers. In  Table 2, we outline the granularity at which different threat models are defined, in the analyzed papers. While most studies examine the exploitability of regular expressions in isolation, some do extend their analysis to the surrounding context, such as within library code. However, only a few studies define threat models at the application level. Threat models defined at code fragment level can lead to false positives—vulnerabilities that may be exploitable in isolated regular expressions but remain unreachable in actual code.
@@
Attacker Controls Input (35 studies): The assumption that an attacker can control inputs to a vulnerable ReDoS (Regular Expression Denial of Service) location is crucial for exploitability, as highlighted in various studies (Barlas et al., 2022; Hassan et al., 2023). In real-world software, including web applications and systems, inputs often originate from untrusted sources, thereby increasing the potential for exploitation (Petsios et al., 2017). However, whether these inputs can reach the vulnerable regular expression depends on the surrounding code. All the reviewed papers, whether their threat models focus on the application (Staicu and Pradel, 2018; Barlas et al., 2022; Wüstholz et al., 2017; Demoulin et al., 2019), library (Bhuiyan et al., 2023; Wang et al., 2023; McLaughlin et al., 2022; Li et al., 2021), or code fragment level (Davis et al., 2017; Turoňová et al., 2022; Liu et al., 2021), assume that the attacker can manipulate the input to exploit the vulnerability in the regular expression. However, this assumption may not always be applicable. For instance, consider a library that only reads and parses a configuration file on a server, which is not controlled by an attacker. In such cases, the likelihood of an attacker being able to influence the input and exploit the ReDoS vulnerability is significantly reduced, if not eliminated.
Attacker’s Input Reaches Super-Linear Regex (34 studies): This condition necessitates that the attacker’s input be matched against a specific type of regular expression. While all the papers considering threat models at the library and code fragment levels implicitly assume this condition, it is a significant assumption. With the exception of Barlas et al. (Barlas et al., 2022), most studies presume this might be achievable in practice. In reality, achieving this could require the attacker to have access to detailed information about the server-side logic, API documentation, or the internal workings of the web service. Staicu et al. (Staicu and Pradel, 2018) incorporate web server implementations into their threat model and examine whether the input can trigger the vulnerability. However, this fingerprinting approach can be noisy, leading to both false positives and false negatives.
@@
Arbitrary Request Assumption (32 studies): This condition represents a strong assumption that the attacker can repeatedly send requests with large input lengths without interference from other security mechanisms. Most studies, with the exceptions of (Shen et al., 2018; Liu et al., 2021; Parolini and Miné, 2023), either disregard restrictions on request size and frequency or necessitate large inputs to trigger vulnerabilities. For context, the maximum size of HTTP header requests and responses in nginx and Apache Tomcat (apa, [n. d.]; ngi, [n. d.]) is 8K, which is significantly smaller than the input lengths assumed in most works. For instance, Davis et al. (Davis et al., 2018a) require input lengths of 100K-1M to trigger a 10-second slowdown, which would be impractical in real-world system settings. Furthermore, most studies fail to consider potential mitigation strategies employed by real-world systems, such as rate limiting(her, [n. d.]), intrusion detection systems, or firewalls, which can significantly impede attacks. In realistic scenarios, attackers might face various system-imposed limitations, including bandwidth constraints or restrictions on the number of requests(her, [n. d.]) they can send. As a result, the majority of studies, except for (Shen et al., 2018; Liu et al., 2021; Su et al., 2024), assume in their threat models that attackers can perform requests unhindered, which may not reflect real-world conditions.

### v4

@@ 3.1.1. Detection
Dynamic Analysis (4 studies): Dynamic approaches (Shen et al., 2018; McLaughlin et al., 2022; Barlas et al., 2022; Petsios et al., 2017) to detecting ReDoS vulnerabilities involve executing regexes with various inputs and analyzing their runtime behavior to identify potential issues. For instance, Shen et al. (Shen et al., 2018) developed ReScue, a tool that uses an evolutionary genetic algorithm to generate time-consuming strings capable of triggering ReDoS vulnerabilities. Similarly, Barlas et al. (Barlas et al., 2022) and McLaughlin et al. (McLaughlin et al., 2022) utilized dynamic analysis with fuzzing techniques, while Petsios et al. (Petsios et al., 2017) introduced SlowFuzz, which uses automated fuzzing to find inputs causing worst-case algorithmic behavior.
Hybrid Approaches (6 studies): Existing static analysis methods for detecting ReDoS vulnerabilities often face a trade-off between precision and recall. For example, Davis et al. (Davis et al., 2018a) and Shen et al. (Shen et al., 2018) reported low F-1 scores of 44.94% and 3.57%, respectively (Li et al., 2021a). To address these limitations, dynamic validation is frequently employed alongside static analysis, resulting in hybrid approaches. Tools such as NFAA (Wüstholz et al., 2017), Revealer (Liu et al., 2021), ReDoSHunter (Li et al., 2021a), Badger (Noller et al., 2018), and RENGAR (Wang et al., 2023) combine static and dynamic methods to mitigate false positives. For instance, hybrid approaches like ReDoSHunter integrate static and dynamic techniques to improve detection accuracy, while tools such as NFAA refine dynamic testing with insights from static analysis (Wüstholz et al., 2017). These methods effectively balance the strengths of both static and dynamic detection, reducing the likelihood of inaccurate results.
Machine Learning (ML)-based Detection (1 study): ML techniques are also gaining traction in ReDoS detection. Unlike hybrid methods, which explicitly integrate static and dynamic analysis, ML approaches aim to predict vulnerabilities and detect anomalies based on patterns in data. For example, Demoulin et al. (Demoulin et al., 2019) introduced a semi-supervised learning model to analyze resource utilization and identify potential vulnerabilities. This method triggers alerts and provides mitigation strategies by detecting anomalies in resource consumption. While ML approaches offer promise, they face challenges such as reliance on labeled data, vulnerability to adversarial inputs, and difficulty in addressing context-specific patterns that may not generalize well across diverse regex use cases.
@@ 3.1.2. Prevention
Better Regex Matching Algorithms (3 studies): Researchers have proposed developing new regex matching engines to address ReDoS vulnerabilities at a fundamental level. Davis et al. (Davis et al., 2021; Davis, 2019) introduced memoization-based optimization to speed up regex matching by caching intermediate results, significantly reducing redundant computations. Turoňová et al. (Turoňová et al., 2020, 2022) proposed novel counting automata matching algorithms that further optimize counter management during matching to efficiently handle regexes with bounded repetitions. These new engines often employ finite-state machine-based approaches instead of backtracking, inherently avoiding the exponential time complexity that leads to ReDoS vulnerabilities. Additionally, some methods impose restrictions on regex features to ensure faster and safer processing.
Automatic Repair (4 studies): Automatic repair techniques aim to proactively address vulnerabilities in regexes to mitigate ReDoS attacks. Several studies have explored programmatic frameworks to detect and repair problematic regex patterns. For example, Li et al. (Li et al., 2021b) developed FlashRegex, a programming-by-example (PbE) tool that repairs or synthesizes regex from provided matching examples. However, FlashRegex struggles with complex regex features, such as lookarounds and backreferences (Li et al., 2022). To address these limitations, RegexScalpel (Li et al., 2022) employs a localize-and-fix approach, which identifies fine-grained vulnerabilities and applies predefined repair strategies to improve recall and precision. In a complementary effort, Hassan et al. (Hassan et al., 2023) focus on identifying specific regex structures or “anti-patterns” that are prone to ReDoS vulnerabilities. Their method analyzes regex patterns to uncover and repair potential vulnerabilities early in the development process, avoiding the need for runtime execution. This approach complements programmatic frameworks by systematically identifying and addressing problematic constructs.
Other tools, such as Chida et al. (Chida and Terauchi, 2022), also leverage PbE algorithms but face challenges related to dependency on user-provided examples, which can result in incomplete repairs (Li et al., 2021a; Wang et al., 2023). These methods, while effective for certain regex patterns, often struggle with advanced or highly customized regex constructs, underscoring the need for more comprehensive solutions capable of addressing complex vulnerabilities. Moreover, improving the generalization and robustness of automatic repair frameworks remains a key area of ongoing research.
@@ 3.1.3. Mitigation
Anomaly Detection (2 studies): Anomaly detection techniques are central to mitigation strategies, leveraging real-time monitoring to identify unusual patterns indicative of ReDoS attacks. ML-based approaches have proven effective in this domain (Bai et al., 2021; Tandon et al., 2023). For example, Tandon et al. (Tandon et al., 2023) employ the Elliptic Envelope model to analyze runtime features such as request time, memory usage, and CPU cycles, enabling the system to detect deviations from normal behavior. Bai et al. (Bai et al., 2021) introduce a feedback loop mechanism that continuously adjusts defensive measures based on real-time performance metrics, providing robust protection against zero-day ReDoS attacks.
@@
Limiting Resource Consumption (1 study): To mitigate ReDoS attacks, Davis et al. (Davis et al., 2018b) proposed using first-class timeouts as a defense against Event Handler Poisoning (EHP) in Node.js. Their prototype, Node.cure, effectively limits processing time with minimal performance overhead. Similarly, runtime environments like PHP, Perl, and .NET have adopted resource-limiting mechanisms to curb backtracking and excessive processing. In PHP, the pcre.backtrack_limit and pcre.recursion_limit directives control the number of backtracks and recursion depth during regex matching (PHP Documentation Group, 2024). Perl’s PCRE2 library allows setting evaluation limits via the eval function, with a default recursion depth limit of 10M. In .NET, the RegexOptions.MatchTimeout property enables developers to specify a maximum regex processing duration. While these strategies effectively cap resources to reduce ReDoS risks, they are often too coarse-grained, making it challenging to balance performance and avoid false positives in complex regex patterns or large datasets.
@@ 3.1.4. Other Related Work
Dataset Studies (1 study): Bhuiyan et al. (Bhuiyan et al., 2023) introduced SecBenchJS, a benchmark dataset for evaluating the security of JavaScript applications, including regex vulnerabilities. Such datasets enable standardized testing and improve ReDoS detection tools.
Empirical Studies (2 studies): Staicu et al. (Staicu and Pradel, 2018) analyzed 2846 popular websites, finding 339 vulnerable to ReDoS attacks, highlighting the risks of server-side JavaScript. Davis et al. (Davis et al., 2019) studied regex portability across eight programming languages, uncovering semantic and performance differences in 25% of regexes. These studies underscore the need for safer regex practices and consistent regex engine implementations.
LLM-based Studies (2 studies): Siddiq et al. (Siddiq et al., 2024b) proposed Re(gEx—DoS)Eval, a framework for evaluating LLM-generated regexes using metrics like pass@k and vulnerable@k, demonstrating its use on T5, Phi-1.5, and GPT-3.5-Turbo. In another study, Siddiq et al. (Siddiq et al., 2024a) categorized vulnerable regex patterns and analyzed developer discussions, finding that GPT-3.5-Turbo often generates regexes with polynomial vulnerabilities.
@@ 3.2. Threat Models
Table 2. Overview of ReDoS threat models in prior work. Each column shows whether the study analyzes code fragments (standalone regex snippets), use within libraries, or full applications, and whether it assumes the attacker can send an arbitrary number of requests. The symbol ● denotes the paper covers that aspect in detail; – means it does not.
Paper Code Fragment Library Application Arbitrary Number of Requests
Su et al. (Su et al., 2024) 100kB
Davis et al. (Davis et al., 2017)
Turoňová et al. (Turoňová et al., 2022) 1.5K–9K
Liu et al. (Liu et al., 2021) 128
Staicu et al. (Staicu and Pradel, 2018) 82K
Davis et al. (Davis et al., 2018a) 100K–1M
Siddiq et al. (Siddiq et al., 2024a)
Shen et al. (Shen et al., 2018) 128
Barlas et al. (Barlas et al., 2022)
Siddiq et al. (Siddiq et al., 2024b)
Davis et al. (Davis et al., 2019)
Kluban et al. (Kluban et al., 2022)
Bhuiyan et al. (Bhuiyan et al., 2023)
Noller et al. (Noller et al., 2018)
Wang et al. (Wang et al., 2023)
Rathnayake et al. (Rathnayake and Thielecke, 2017)
McLaughlin et al. (McLaughlin et al., 2022) 1M
Li et al. (Li et al., 2021a)
Parolini et al. (Parolini and Miné, 2023) 128
Kirrage et al. (Kirrage et al., 2013)
Wüstholz et al. (Wüstholz et al., 2017) 140K
Petsios et al. (Petsios et al., 2017)
Kluban et al. (Kluban et al., 2024)
Demoulin et al. (Demoulin et al., 2019)
Chida et al. (Chida and Terauchi, 2022)
Bai et al. (Bai et al., 2021) 50K
Li et al. (Li et al., 2022) 1M
Davis et al. (Davis, 2019)
Li et al. (Li et al., 2021b)
Hassan et al. (Hassan et al., 2023)
Davis et al. (Davis et al., 2018b)
Davis et al. (Davis et al., 2021) 10K–100K
Tandon et al. (Tandon et al., 2023)
Weideman et al. (Weideman et al., 2016)
Turoňová et al. (Turoňová et al., 2020) 500K–50M
A realistic threat model is essential in any security study, as it provides a framework for assessing the potential findings. That is, it enables the identification of the system’s weaknesses (i.e., vulnerabilities) that adversaries could exploit in real-world environments. A strong threat model might lead to unrealistic findings, assuming too powerful attackers. In  Table 2, we outline the granularity at which different threat models are defined in the analyzed papers. While most studies examine the exploitability of regexes in isolation, some do extend their analysis to the surrounding context, such as within library code. However, only a few studies define threat models at the application level. Threat models defined at the code fragment level can lead to false positives—vulnerabilities that are exploitable in isolated regexes but unreachable in actual code.
@@
Attacker Controls Input (35 studies): The assumption that an attacker can control inputs to a vulnerable ReDoS location is crucial for exploitability, as highlighted in various studies (Barlas et al., 2022; Hassan et al., 2023). In real-world software, including web applications and systems, inputs often originate from untrusted sources, thereby increasing the potential for exploitation (Petsios et al., 2017). However, whether these inputs can reach the vulnerable regex depends on the surrounding code. All the reviewed papers, whether their threat models focus on the application (Staicu and Pradel, 2018; Barlas et al., 2022; Wüstholz et al., 2017; Demoulin et al., 2019), library (Bhuiyan et al., 2023; Wang et al., 2023; McLaughlin et al., 2022; Li et al., 2021a), or code fragment level (Davis et al., 2017; Turoňová et al., 2022; Liu et al., 2021), assume that the attacker can manipulate the input to exploit the vulnerability in the regex. However, this assumption may not always be applicable. For instance, consider a library that only reads and parses a configuration file on a server, which is not controlled by an attacker. In such cases, the likelihood of an attacker being able to influence the input and exploit the ReDoS vulnerability is significantly reduced, if not eliminated.
Attacker’s Input Reaches Super-Linear Regex (34 studies): This condition necessitates that the attacker’s input be matched against a specific type of regex. While all the papers considering threat models at the library and code fragment levels implicitly assume this condition, it is a significant assumption. With the exception of Barlas et al. (Barlas et al., 2022), most studies presume this might be achievable in practice. In reality, achieving this could require the attacker to have access to detailed information about the server-side logic, API documentation, or the internal workings of the web service. Staicu et al. (Staicu and Pradel, 2018) incorporate web server implementations into their threat model and examine whether the input can trigger the vulnerability. However, this fingerprinting approach can be noisy, leading to both false positives and false negatives.
@@
Arbitrary Request Assumption (32 studies): This condition represents a strong assumption that the attacker can repeatedly send requests with large input lengths without interference from other security mechanisms. Most studies, with the exceptions of (Shen et al., 2018; Liu et al., 2021; Parolini and Miné, 2023), either disregard restrictions on request size and frequency or necessitate large inputs to trigger vulnerabilities. For context, the maximum size of HTTP header requests and responses in nginx and Apache Tomcat (Foundation, 2018; Community, 2024a) is 8K, which is significantly smaller than the input lengths assumed in most works. For instance, Davis et al. (Davis et al., 2018a) require input lengths of 100K-1M to trigger a 10-second slowdown, which would be impractical in real-world system settings. Furthermore, most studies fail to consider potential mitigation strategies employed by real-world systems, such as rate limiting, intrusion detection systems, or firewalls, which can significantly impede attacks. In realistic scenarios, attackers might face various system-imposed limitations, including bandwidth constraints or restrictions on the number of requests (Salesforce, 2024) they can send. As a result, the majority of studies, except for (Shen et al., 2018; Liu et al., 2021; Su et al., 2024), assume in their threat models that attackers can perform requests unhindered, which may not reflect real-world conditions.